O 4/30/2016 11:55 PM A 33.3/100 Gradebook Print Ca
Answer #1

12H + 36Li ---> 37Li + 11H

Total mass number A,

RHS=2+6=8

Total atomic number

RHS=1+3=4

So,

That unknown element have the mass number of 7 and atomic number of 3.

Hence, 37Li is the unknown element.

Sum of the masses in RHS of the reaction,

The exact mass of 12H = 2.01410 a.m.u.

3Li6 =6.015123 a.m.u.

Total = 8.029223 a.m.u.

LHS

The exact mass of 3Li7 isotope = 7.01601 a.m.u.

1H1 = 1.00738 a.m.u.

Total = 8.02384 a.m.u.

ΔM = Sum of the masses of products – Sum of the masses of reactants

= 8.02384 -  8.029223

= - 0.005383 a.m.u.

As there is decrease of mass, hence energy is being released.

The value of Q will be therefore Positive i.e.
Q = +0.005383 a.m.u. X 931.5 MeV/a.m.u = +5.011573 MeV
Hence the above reaction is exoergic.